OpenStudy (anonymous):

simplify: (3√8)-(7√18)

OpenStudy (anonymous):

ok 8= 2³ and 18= 3²x2so (3√8)-(7√18)=6√2-21√2

OpenStudy (anonymous):

\[3\sqrt{8}=3\sqrt{2*4}=6\sqrt{2}\] \[7\sqrt{18}=7\sqrt{9*2}=21\sqrt{2}\]

OpenStudy (anonymous):

p-(q\p-q)-q-(q-p) =??